40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10960     Entry time: Fri Jan 30 03:12:15 2015
Author: diego 
Type: Update 
Category: LSC 
Subject: CARM on REFL11I 

[Jenne, Diego]

Tonight we continued following the plan of last night: perform the transition of CARM to REFL11_I while on MICH offset at -25%:

  • we managed to do the transition several times, keeping the UGF servos on for MICH and PRCL but turning off the DARM and CARM ones, because their contribution was rather unimportant and we feared that their excitations could affect negatively the other loops (as loops tend to see each other's excitation lines);
  • we had to tweak the MICH and PRCL UGF servos:
    • the excitation frequency for MICH was lowered to ~41 Hz, while PRCL's one was lowered to ~50 Hz;
    • PRCL's amplitude was lowered to 75 because it was probably too high and it affected the CARM loop, while MICH's one was increased to 300 because during the reduction of the CARM offset it was sinking into the noise; after a few tries we can say they don't need to be tweaked on the fly during the procedure but can be kept fixed from the beginning;
    • after the transition to REFL11_I for CARM, we engaged also its UGF servo, still at the highest frequency of the lot (~115 Hz) and with relatively low amplitude (2), to help keeping the loop stable;
    • as DARM was still on ALS, we didn't engage its UGF servo during or after the transition, but we just hold its output from the initial part of the locking sequence (after we lowered its frequency to 100 Hz;
  • however, at CARM offset 0 our arm power was less that what we had yesterday: we managed to get higher than ~8, but after Koji tweaked the MC alignment we reached ~10; we still don't understand the reason of the big difference with respect to what the simulations show for MICH offset at 25% (arm power ~50);
  • after the CARM transition to REFL11_I we felt things were pretty stable, so we tried to reduce the MICH offset to get us in the ~ -10% range, however we never managed to get past ~ -15% before losing lock, at arm power around 20;
  • we lost lock several times, but for several different reasons (IMC lost lock a couple of times, PRCL noise increased/showed some ringing, MICH railed) but our main concern is with the PRCL loop:
    • we took several measurements of the PRCL loop: the first one seemed pretty good, and it had a bigger phase bubble than usual; however, the subsequent measurements showed some weird shapes we struggle to find a reason for; these measurements were taken at different UGF frequencies, so maybe it is worth looking for some kind of correlation; morever, in the two weird measurements the UGFs are not where they are supposed to be, even if the servo was correctly following the input (or so it seemed); the last measurement was interrupted just before we lost lock because of PRCL itself;
    • we noticed a few times during the night that the PRCL loop noise in the 300-500 Hz range increased suddenly and we saw some ringing; at least a couple of times it was PRCL who threw us out of lock; this frequency range is similar to the 'weird' range we found in our measurements, so we definitely need to keep an eye on PRCL on those frequencies;
  • in conclusion, the farthest we got tonight was CARM on REFL11_I at 0 offset, DARM at 0 offset still on ALS and MICH at ~ 15% offset, arm power ~20.

 

Attachment 1: PRCL_29Jan2015_Weird_Shape.pdf  7 kB  | Hide | Hide all
PRCL_29Jan2015_Weird_Shape.pdf
Attachment 2: ArmPowers20_MICHoffsetBeingReduced_0CARMoffset_29Jan2015.pdf  28 kB  Uploaded Fri Jan 30 04:12:31 2015  | Hide | Hide all
ArmPowers20_MICHoffsetBeingReduced_0CARMoffset_29Jan2015.pdf
ELOG V3.1.3-